OpenStudy (anonymous):

The tables below show the values of f(x) and g(x) for different values of x: f(x) = 2(3)x x f(x) −2 0.22 −1 0.67 0 2 1 6 2 18 g(x) = 3x + 3 x g(x) −2 3.11 −1 3.33 0 4 1 6 2 12 Based on the tables, what is the solution to the equation 2(3^)x = 3^x + 3? A. x = −1 B. x = 0 C. x = 1 D. x = 6

OpenStudy (mathmate):

Examine both tables. The solution is where the two functions f(x) and g(x) intersect. When two graphs (functions) intersect, they have identical x,y values. Therefore, look for identical rows between the two functions, i.e. where both x and y are identical. The question requires only the value of x in the x-y pair.

OpenStudy (anonymous):

i don't get it @mathmate

OpenStudy (mathmate):

@Nicole_gu Look at the first line of f(x), what do you see?

OpenStudy (anonymous):

is it c??

OpenStudy (mathmate):

By the way, the equations should have been written: f(x)=2(3)^x, and g(x)=3^x+3 When you see an exponent, you need to show others that it is an exponent (with the ^, or the caret).

OpenStudy (mathmate):

Would you like to tell me what you see in the first lines of f(x) and g(x)?

OpenStudy (anonymous):

1

OpenStudy (mathmate):

Can you type out the whole line that you see? The first line does not show a "1".

OpenStudy (anonymous):

i don't know its that i really have to go so can you explain it to me please i have to go!!

OpenStudy (mathmate):

The first line for f(x) is -2 0.22 and the first line for g(x) is -2 3.11 ok, so we'll continue later! Bye! :)